OpenStudy (anonymous):

2x+5y=2 -3x-y=-3 im supposed to solve by substitution and im stuck

OpenStudy (chosenmatt):

what numbers will make the equation true?

OpenStudy (chosenmatt):

what ever number you choose for x times that number by the number beside x

OpenStudy (anonymous):

i dont know all i know is that the answer is supposed to be in (x,y) form

OpenStudy (chosenmatt):

same for y

OpenStudy (chosenmatt):

@ant501

OpenStudy (anonymous):

thankyou for helping @chosenmatt

OpenStudy (anonymous):

what matt

OpenStudy (chosenmatt):

@HelpBlahBlahBlah am i right?

OpenStudy (anonymous):

wait a sec gotta fix I gotta fix this again

OpenStudy (anonymous):

back

OpenStudy (anonymous):

lol i dont know really i suck at algebra but you might be and even if you arent im sure your closer to the answer than me i just started solving substition problems

OpenStudy (anonymous):

i just appreciate that you took the time to try to help

OpenStudy (kendricklamar2014):

Step 1: Simplify both sides of the equation. 2x−15=−3x+5

OpenStudy (kendricklamar2014):

Step 2: Add 3x to both sides. 2x−15+3x=−3x+5+3x 5x−15=5

OpenStudy (kendricklamar2014):

Step 3: Divide both sides by 2. 2x/2 = 17/2

OpenStudy (kendricklamar2014):

X = 17/2